# Robust direction-dependent gain-calibration of beam-modelling errors far from the target field

### v1

@@ 1 Introduction
However, knowledge of the on-sky sources alone is not sufficient to properly remove them from the data, because propagation and instrumental effects affect their signature in the data. For example, ionospheric effects can cause the sources to scintillate (Cronyn, 1972; Koopmans, 2010), and the chromatic components of the instrument itself may cause imprints on all data (de Gasperin et al., 2019). These effects are corrected for during calibration. For wide fields of view, the most important are the temporally and spatially varying ionospheric and station-beam effects (Lonsdale, 2005; Smirnov, 2011b). The two main approaches to removing these are demixing (van der Tol et al., 2007) and direction-dependent (DD) calibration. Although demixing (i.e. self-calibration on visibilities that have been phase-rotated to the source that needs to be removed) can be effective for removing very bright off-axis sources, it is not as versatile as DD-gain calibration. Demixing is only able to calibrate for a single direction at a time, such that such a source must dominate the field around it. DD-gain calibration, on the other hand, can cover a larger range of signal-to-noise ratios and source separations. In DD-gain calibration, different gains are calculated in the directions of groups of sources, rather than for the full sky. This allows for the necessary flexibility to account for DD-gain errors, but increases the number of degrees of freedom in the calibration process, which creates a risk of overfitting in the gain solutions (Patil et al., 2016; Ewall-Wice et al., 2017; Mouri Sardarabadi & Koopmans, 2019). Such an overfit can lead to suppression of the signal of interest. Therefore, additional constraints have to be introduced to regularise the calibration solutions. Because both propagation effects and individual receiver gains are physically expected to be smooth over a large range of frequencies, one of the most commonly imposed constraints is spectral smoothness. Spectral smoothness is a permissible assumption for most ionospheric effects, but can falter in areas where the instrument primary beam is highly spectrally irregular (Trott & Wayth, 2016), due to the phasing of several receivers to form a tile or station beam.
@@
Another possibility is to calibrate with an improved model for the beam, such that the most rapid variations are already built into the beam model rather than absorbed into the gains (e.g. Line et al. 2018; Iheanetu et al. 2019; Nunhokee et al. 2020). This is challenging for phased arrays, however, as their beams change as a function of pointing direction, such that models with many parameters are needed to describe all possible pointings. Fully numerical simulations are not feasible to the required accuracy, because the in situ behaviour of the individual antenna elements is difficult to predict (e.g. Wilensky et al. 2024). Also, failures or degradations within the station may occur. These are not straightforward to detect post-commissioning (Chokshi et al., 2024). Another option for creating improved beam models is holography using a bright source. One can introduce an artificial bright source (e.g. drone). This is often impossible, however, both because of the many holographic measurements that would be needed to cover the range of possible pointing directions, and because of the distance to the far field. This far field can start kilometres up in the air due to the station sizes, such that flying a sufficiently bright calibrator source over the instrument is not feasible (Jacobs et al., 2017). Using satellite carrier emission or reflections instead is a possibility, but these signals are typically narrow-band and therefore cannot be used to create an accurate model across the full spectral band of the instrument. Finally, bright sky-sources can be used, but only in a limited set of directions. To create a model across the full band and for all pointings, most modern radio observatories opt for simplified beam models instead. However, these falter in the presence of real-world effects, such as broken antenna elements, drifting or degraded amplifiers, or electromagnetic coupling within the station. Although these models are useful to first order and sufficient in many cases, they do leave errors which are prohibitive in the most ambitious science cases.
@@
We illustrate the effect of DD beam-errors on gain-calibration results using the EoR science case and the LOFAR-HBA instrument, because this example is especially affected by DD errors (Patil et al., 2016; Ewall-Wice et al., 2017; Sardarabadi & Koopmans, 2019). These errors are likely one of the dominant causes of unwanted residual emission ("excess variance") seen in the data after sky model subtraction with the DD-gain solutions applied (Gan et al., 2022; Mertens et al., 2025). Beam errors have also been shown to be limiting this science case on similar instruments such as the MWA (Chokshi et al., 2024) and NenuFAR (Munshi et al., 2024). Because the short baselines that are the main focus of interferometric EoR science are not strongly affected by the ionosphere (Brackenhoff et al., 2024), we can focus mainly on the beam itself.
We present the impact of realistic beam errors on LOFAR-HBA by switching off individual antenna elements within the station, as this is expected to be a dominant source of beam modelling errors and differences between individual LOFAR-HBA stations. We calculate the resulting beam changes, the effects they have on the visibility data, and the impact on the spectrally regularised calibration of sources far from the target direction. This is the first simulation that shows the full end-to-end effect of beam errors on LOFAR-HBA DD-calibration on the scale of a full night of observations. Furthermore, we present a first potential heuristic solution to the broken spectral smoothness assumption in DD-gain calibration and show that this solution improves upon the standard method both in visibility and image space. In this more robust regularisation method, the regulariser is weighted by the forward prediction of the visibility contribution of each direction, such that time- and frequency intervals with strong beam attenuation do not negatively affect the gain solutions at other times and frequencies. We use the observation and calibration settings from LOFAR EoR to illustrate the impact of beam errors, but emphasize that these results apply to many science cases and instruments, especially those needing the shorter baselines.
@@
Notation: Throughout this work, we indicate vectors with lowercase bold italic typesetting ($\bm{a}$), unit-length vectors with a hat ($\bm{\hat{a}}$) and matrices with uppercase bold sans-serif typesetting ($\bm{\mathsf{A}}$). The complex conjugate, complex conjugate transpose and matrix inverse are indicated by $(\cdot)^{*}$, $(\cdot)^{H}$, and $(\cdot)^{-1}$, respectively. The Frobenius norm (the square root of the sum of squares of the elements of a matrix) is denoted by $\|\cdot\|_{F}$. The inner product is denoted with $\cdot$, the Hadamard product by $\odot$ and the imaginary unit by $i$.
2 Direction-dependent gain-calibration outisde the main beam lobe
Radio-interferometric measurements are created by correlating measured voltages between stations. Such a measured correlation, averaged over a small spectrotemporal interval, forms a visibility. Visibilities therefore correspond to a pair of stations [^2] They may also be the auto-correlation of a station with itself, that form a baseline. Using the Jones formalism, one can write the visibility on the baseline between stations $p$ and $q$ at time $t$ and frequency $\nu$ with a $2\times 2$ matrix $\bm{\mathsf{V}}_{p,q}(\nu,t)$. This matrix is often described using the Radio Interferometric Measurement Equation (RIME, Hamaker et al. 1996; Smirnov 2011a). For a single point-source, the RIME is
@@ 2 Direction-dependent gain-calibration outisde the main beam lobe
In this equation, $\bm{\mathsf{C}}_{p,q}$ is the $2\times 2$ source brightness coherency matrix, which describes the signal from the sky. $\bm{\mathsf{P}}_{p}$ is a $2\times 2$ propagation matrix describing the signal path of incident radiation on station $p$. Finally, $\bm{\mathsf{N}}_{p,q}$ consists of white Gaussian noise [^3] Note that the elements of $\bm{\mathsf{N}}$ describe different polarizations and not the statistical behaviour of the noise.
For uncorrelated, discrete sources, the brightness of the full sky is additive. If we assume all components are sufficiently compact that the Jones matrix can be assumed constant over the extent of the source, the full-sky RIME can be broken up into a set of components denoted by $d$, i.e.
@@
A phased array can determine the angle of arrival from the delay by which a wavefront is received by its elements. A wavefront originating from a source in direction $\bm{\hat{s}}$ has a wave vector of $\bm{k}(\bm{\hat{s}},\nu)=-\frac{2\pi\nu}{c}\bm{\hat{s}}$ [m-1], where $c$ is the speed of light in metres per second. For a wavefront received by a phased array station $p$ consisting of $N_{m}$ elements at positions $\bm{x}_{p,m}$, these delays can be written as $\exp\left(-i\bm{k}\cdot\bm{x}_{p,m}\right)$. Therefore, the station response to a source can be modelled as
@@
where $\bm{\mathsf{G}}_{p,m}$ are per-element complex gains, $\bm{\hat{s}_{0}}$ is the phase centre, and $\bm{\mathsf{U}}_{p}(\bm{\hat{s}},\nu)$ is the element beam, which we assume to be identical between elements here.
The propagation matrix is then split into this beam model, and a model for all other propagation effects (both astronomical and instrumental), i.e. $\bm{\mathsf{P}}_{p,d}=\bm{\mathsf{J}}_{p,d}\bm{\mathsf{A}}_{p,d}$. Calibration then attempts to find the matrices $\bm{\mathsf{J}}_{p,d}(\nu,t)$ that make this sky model best match the data. Because the sky usually remains coherent over short time and frequency intervals, a matrix $\bm{\mathsf{J}}_{p}$ is found over such an interval instead of per data point, both raising the signal-to-noise ratio and reducing the computational cost. We define a parameter $l$ that iterates over all combinations of $\nu$ and $t$ within a calibration interval to simplify notation. For similar reasons, groups of spatially close sources (calibration clusters, denoted by $k$) are assigned the same Jones matrix $\bm{\mathsf{J}}_{p,k}$, rather than an individual matrix per source $\bm{\mathsf{J}}_{p,d}$. This is an approximation, since the clustering of sources is a heuristic procedure based on the required minimum flux level of a calibration cluster, or the total permissible number of calibration clusters. This means that one pair of closely spaced sources can be assigned the same calibration cluster, whereas a similarly spaced source is assigned a different matrix, even though its propagation path can be very similar to that of the original pair. If the clusters are not well-chosen, the gain can abruptly change on the boundaries between calibration clusters, effectively imposing a top-hat function on the spatial variations in propagation. The matrix $\bm{\mathsf{J}}_{p,k}$ models the typical propagation effects for sources in a calibration cluster, but is an approximation for each source individually. The simplest form of the calibration problem then becomes
$\underset{\bm{\mathsf{J}}_{p,k}}{\text{Minimise }}$ $\sum_{p,q}\ \sum_{l}$
@@
Any imperfections in the (lack of) beam model are absorbed into the calculated solutions $\bm{\mathsf{J}}_{p,k}$.
@@ 2.1 DDECal
DDECAL contains several algorithms [^6] These are directionsolve, directioniterative, hybrid and LBFGS. The LBFGS routine was implemented in DDECAL specifically to compare results to those obtained with SAGECAL-CO and not widely used in isolation. directioniterative and hybrid are mainly used to reduce the required computational time, which is why we focus on the more precise directionsolve in this work. that numerically approximate the solution of Equation 4. Below, we describe its directionsolve algorithm, which is similar to the algorithm described by Smirnov & Tasse (2015), but with the addition of a constraints framework. The solver finds the gains by solving for one station in all directions simultaneously. For all other stations, the gains of the previous iteration are assumed. In Equation 4, this means that $\bm{\mathsf{J}}_{q,k}$ are assumed to be known, while for each $p$, $\bm{\mathsf{J}}_{p,k}$ are updated for all directions $k$ simultaneously. For each interval, the data and gains are stacked column-wise, as [^7] The notation used here is a corrected version of that presented by Gan et al. (2023), but is consistent with a new version of their notation available at https://arxiv.org/abs/2209.07854v2.,
@@
    where we have dropped the subscript ‘sol’ for legibility. The Gaussian kernel is given by $K(\nu,\nu^{\prime})=\exp[-\frac{1}{2}(\nu-\nu^{\prime})^{2}/\sigma^{2}]$, and $W_{p,k}(\nu)$ is a weight. Such weights are useful when some intervals contain fewer data points than others, due to interference excision, for example.

### v2

@@ 1 Introduction
However, knowledge of the on-sky sources alone is not sufficient to properly remove them from the data, because propagation and instrumental effects affect their signature in the data. For example, ionospheric effects can cause the sources to scintillate (Cronyn, 1972; Koopmans, 2010). Furthermore, the chromatic components of the instrument itself, such as the bandpass filter in LOFAR, may cause imprints on all data (de Gasperin et al., 2019). Calibration attempts to correct for these effects. For wide fields of view, the most important are the temporally and spatially varying ionospheric and primary beam effects (Lonsdale, 2005; Smirnov, 2011b). The two main approaches to removing these are demixing (van der Tol et al., 2007) and direction-dependent (DD) calibration. Although demixing (i.e. self-calibration on visibilities that have been phase-rotated to the source that needs to be removed) can be effective for removing very bright off-axis sources, it is not as versatile as DD-gain calibration. Demixing is only able to calibrate for a single direction at a time, and the source must dominate the field around it. This places constraints on the minimum source brightness and the minimum separation from other (bright) sources in the vicinity. DD-gain calibration, on the other hand, can cover a larger range of signal-to-noise ratios and source separations, but is computationally more expensive. In DD-gain calibration, different gains are calculated in the directions of groups of sources, rather than for the full sky. This allows for the necessary flexibility to account for DD errors, but also increases the number of degrees of freedom in the calibration process, which creates a risk of overfitting in the gain solutions (Patil et al., 2016; Ewall-Wice et al., 2017; Mouri Sardarabadi & Koopmans, 2019). Such an overfit can lead to suppression of the signal of interest. Therefore, additional constraints have to be introduced to regularise the calibration solutions. Because both propagation effects and individual receiver gains are physically expected to be smooth over a large range of frequencies, one of the most commonly imposed constraints is spectral smoothness. Spectral smoothness is a permissible assumption for most ionospheric effects, but can falter in areas where the instrument primary beam is highly spectrally irregular (Trott & Wayth, 2016). Such spectral irregularities occur in phased array stations that consist of several antenna elements that are coherently summed to create a larger station.
@@
Another possibility is to calibrate with an improved model for the beam, such that the most rapid variations are already built into the beam model rather than absorbed into the gains (e.g. Line et al. 2018; Iheanetu et al. 2019; Nunhokee et al. 2020). This is challenging for phased arrays, however, as their beams change as a function of pointing direction, such that models with many parameters are needed to describe all possible pointings. Most interferometric beam models used in calibration therefore assume identical elements. The station beam is computed by using a full electromagnetic simulation of an element beam, and imposing the station structure through a sum of progressive phase shifts (the ‘array factor’), ignoring variations in mutual coupling between antenna elements in the array (Wijnholds et al., 2019). This type of model describes the general behaviour of the beam well, but suffers from possible instability. If the station has a regular structure, a beam model created with identical elements has deep nulls and strong sidelobes, that may shift or disappear when the station symmetry is broken. This can occur when there are variations in the element beam patterns (e.g. due to mutual coupling), when there are gain variations between elements, or when elements break. Broken symmetries have a smoothing effect on the beam pattern. When the beam model used in calibration is less smooth than the true beam, calibration can imprint the structure of the model beam onto the data.
Models that do incorporate mutual coupling effects are under construction (O’Hara et al., 2024), but full numerical electromagnetic simulations are not feasible to the required accuracy, because the in situ behaviour of the individual antenna elements is difficult to predict (e.g. Wilensky et al. 2024). Also, failures or degradations within the station may occur. These are not straightforward to detect post-commissioning (Chokshi et al., 2024). Another option for creating improved beam models is holography using a bright source. One can introduce an artificial bright source (e.g. drone). This is often impossible, however, both because of the many holographic measurements that would be needed to cover the range of possible pointing directions, and because of the distance to the far field. This far field can start kilometres up in the air due to the station sizes, such that flying a sufficiently bright calibrator source over the instrument is not feasible (Jacobs et al., 2017). Using satellite carrier emission or reflections instead is a possibility, but these signals are typically narrow-band and therefore cannot be used to create an accurate model across the full spectral band of the instrument. Finally, bright sky-sources can be used, but only in a limited set of directions (Nunhokee et al., 2020). To create a model across the full band and for all pointings, most modern radio observatories opt for simplified beam models instead. However, these falter in the presence of real-world effects, such as broken antenna elements, drifting or degraded amplifiers, or electromagnetic coupling within the station. Although these models are useful to first order and sufficient in many cases, they do leave errors which are prohibitive in the most ambitious science cases.
@@
We illustrate the effect of DD beam-modelling errors on the gains recovered during DD gain-calibration using the EoR science case and the LOFAR-HBA instrument, because this example is especially affected by DD errors (Patil et al., 2016; Ewall-Wice et al., 2017; Sardarabadi & Koopmans, 2019). These errors are likely one of the dominant causes of unwanted residual emission ("excess variance") seen in the data after sky model subtraction with the DD-gain solutions applied (Gan et al., 2022; Mertens et al., 2025). Beam errors have also been shown to be limiting this science case on similar instruments such as the MWA (Chokshi et al., 2024) and NenuFAR (Munshi et al., 2024). Because the short baselines that are the main focus of interferometric EoR science are not strongly affected by the ionosphere (Brackenhoff et al., 2024), we can focus mainly on the beam itself.
We present the impact of realistic beam errors on LOFAR-HBA with a Python code named SHIMMERR (Station Heterogeneity Impact on Multi-dimensional beam-Modelling Errors simulatoR and calibratoR) [^2] https://github.com/Stefanie-B/shimmerr. We create beam errors by simulating that a number of individual antenna elements within the station are switched off, as this is expected to be a dominant source of beam modelling errors and differences between individual LOFAR-HBA stations. We calculate the resulting beam changes, the effects they have on the visibility data, and the impact on the spectrally regularised calibration of sources far from the target direction. This is the first simulation that shows the full end-to-end effect of beam errors on LOFAR-HBA DD-calibration on the scale of a full night of observations. Furthermore, we present a first potential heuristic solution to the broken spectral smoothness assumption in DD-gain calibration and show that this solution improves upon the standard method both in visibility and image space. In this more robust regularisation method, the regulariser is weighted by the forward prediction of the visibility contribution of each direction, such that time- and frequency intervals with strong beam attenuation do not negatively affect the gain solutions at other times and frequencies. We use the observation and calibration settings from LOFAR EoR to illustrate the impact of beam errors, but emphasize that these results apply to many science cases and instruments, especially those needing the shorter baselines.
@@
[…]
2 Direction-dependent gain-calibration outside the main beam lobe
In radio astronomy, the sky is described by its brightness distribution, denoted by $I(l,m)$, where $l$ and $m$ are direction cosines. $I(l,m)$ represents the specific intensity as a function of sky direction. Under the assumptions of spatial incoherence of $I(l,m)$ and quasi-monochromatic radiation, the van Cittert-Zernike theorem provides the theoretical basis for radio interferometry. The theorem states that the spatial coherence of the electric field (i.e. a visibility) is given by the Fourier transform of the sky brightness distribution $I$ (Thompson et al., 2001). A baseline (i.e. the vector connecting two receiver elements) with components $(u,v,w)$ in units of wavelength measures a visibility $V(u,v,w)$,
[…]
Here, $l,m$ are direction cosines on the sky, and $A(l,m)$ is the beam response of the receiver.
Whilst the van Cittert-Zernike theorem is an excellent tool for understanding the physical meaning of visibilities, it has its limitations, because it can only describe scalar fields (i.e. no polarisation effects), and does not include instrumental or propagation effects. The Van-Cittert Zernike theorem can be generalised to the Radio Interferometric Measurement Equation (RIME, Hamaker et al. 1996; Smirnov 2011a). For a single point-source, the RIME is given by
@@ 2 Direction-dependent gain-calibration outside the main beam lobe
In this equation, $\bm{\mathsf{V}}_{p,q}$ is a $2\times 2$ matrix that describes the visibility on the baseline between receiver elements $p$ and $q$ at time $t$ and frequency $\nu$ with the Jones formalism. $\bm{\mathsf{P}}_{p}$ is a $2\times 2$ propagation matrix describing the signal path of incident radiation on receiver element $p$, as well as instrumental effects. Finally, $\bm{\mathsf{N}}_{p,q}$ describes white Gaussian noise [^3] Note that the elements of $\bm{\mathsf{N}}$ describe different polarisations and not the statistical behaviour of the noise. $\bm{\mathsf{C}}_{p,q}$ is the $2\times 2$ source coherency matrix, which describes the correlations between orthogonal components of the electric field originating from the source. Let $\bm{e}=\left[e_{x},e_{y}\right]^{T}$ be the complex Jones vector that represent the complex electric field of the source. The source coherency matrix is then defined as
$\bm{\mathsf{C}}_{p,q}(\nu,t)=\left<\bm{e}\bm{e}^{H}\right>,$ (3)
[…]
For uncorrelated, discrete sources, the brightness of the full sky is additive, because the electric fields from incoherent sources sum in power rather than amplitude. If we assume all components are sufficiently compact that the Jones matrix can be assumed constant over the extent of the source, the full-sky RIME can be broken up into a set of components denoted by $d$, i.e.
@@
In standard radio interferometry, individual antennas or dishes are used as receiver elements (i.e. directly correlated to form visibilities). However, the receiver elements in many modern interferometers, such as LOFAR, SKA-Low, NenuFAR, and MWA, are phased array stations, rather than single antennas. Each station consists of a set of $N_{m}$ antennas that are coherently summed to create electronically steerable beams. By introducing phase shifts based on the antenna element positions $\bm{x}_{p,m}$ in metres, the station can be steered in direction $\bm{\hat{s}_{0}}$,
[…]
A phased array can determine the angle of arrival from the delay by which a wavefront is received by its elements. A wavefront originating from a source in direction $\bm{\hat{s}}$ has a wave vector of $\bm{k}(\bm{\hat{s}},\nu)=-\frac{2\pi\nu}{c}\bm{\hat{s}}$ [m-1]. These delays can be written as $\exp\left(-i\bm{k}\cdot\bm{x}_{p,m}\right)$. The station response to a source can be modelled as
@@
where $\bm{\mathsf{G}}_{p,m}$ are per-element complex gains, and $\bm{\mathsf{U}}_{p}(\bm{\hat{s}},\nu)$ is the antenna element beam (e.g. a simple dipole), which we assume to be identical between antennas here. The element-beam independent part of this equation,
$\mathrm{AF}(\bm{\hat{s}},\nu)=\sum_{m}\bm{\mathsf{G}}_{p,m}\exp\left[-i\bm{k}(\bm{\hat{s}}-\bm{\hat{s}_{0}},\nu)\cdot\bm{x}_{p,m}\right],$ (8)
is called the ‘array factor’. It represents the far-field beam of the station based solely on the element positions and their excitations, assuming identical, isotropic elements.
The propagation matrix is then split into the beam model, and a model for all other propagation effects (both astronomical and instrumental), i.e. $\bm{\mathsf{P}}_{p,d}=\bm{\mathsf{J}}_{p,d}\bm{\mathsf{A}}_{p,d}$. Calibration then attempts to find the matrices $\bm{\mathsf{J}}_{p,d}(\nu,t)$ that make this sky model best match the data. Because the sky usually remains coherent over short time and frequency intervals, a matrix $\bm{\mathsf{J}}_{p}$ is found over such an interval instead of per data point, both raising the signal-to-noise ratio and reducing the computational cost. We call these intervals ‘solution intervals’ and define a parameter $l$ that iterates over all combinations of $\nu$ and $t$ within a solution interval to simplify notation. For similar reasons, groups of spatially close sources (calibration clusters, denoted by $k$) are assigned the same Jones matrix $\bm{\mathsf{J}}_{p,k}$, rather than an individual matrix per source $\bm{\mathsf{J}}_{p,d}$. This is an approximation, since the clustering of sources is a heuristic procedure based on the required minimum flux level of a calibration cluster, or the total permissible number of calibration clusters. This means that one pair of closely spaced sources can be assigned the same calibration cluster, whereas another similarly spaced source could be assigned a different matrix, even though its propagation path can be very similar to that of the original pair. If the clusters are not well-chosen, the gain can abruptly change on the boundaries between calibration clusters, effectively imposing a two-dimensional step-function on the spatial variations in propagation. The matrix $\bm{\mathsf{J}}_{p,k}$ models the typical propagation effects for sources in a calibration cluster, but is an approximation for each source individually. The simplest form of the calibration problem, for each solution interval separately, then becomes
$\underset{\bm{\mathsf{J}}_{p,k}}{\text{Minimise }}$ $\sum_{p,q}\ \sum_{l}$ (9)
@@
The aim is then to absorb the largest imperfections in the (lack of) beam model into the calculated solutions $\bm{\mathsf{J}}_{p,k}$.
@@ 2.1 DDECal
DDECAL contains several algorithms [^6] These are directionsolve, directioniterative, hybrid and LBFGS. The LBFGS routine was implemented in DDECAL specifically to compare results to those obtained with SAGECAL-CO and not widely used in isolation. directioniterative and hybrid are mainly used to reduce the required computational time, which is why we focus on the more precise directionsolve in this work. that numerically approximate the solution of Equation 10. Below, we describe its directionsolve algorithm, which is similar to the algorithm described by Smirnov & Tasse (2015), but with the addition of a constraints framework. The solver finds the gains by solving for one station in all directions simultaneously. For all other stations, the gains of the previous iteration are assumed. In Equation 10, this means that $\bm{\mathsf{J}}_{q,k}$ are assumed to be known, while for each $p$, $\bm{\mathsf{J}}_{p,k}$ are updated for all directions $k$ simultaneously. For each solution interval, the data and gains are stacked column-wise, as [^7] The notation used here is a corrected version of that presented by Gan et al. (2023), but is consistent with a new version of their notation available at https://arxiv.org/abs/2209.07854v2.,
@@
    where we have dropped the subscript ‘sol’ for legibility. The Gaussian kernel is given by $K(\nu,\nu^{\prime})=\exp[-\frac{1}{2}(\nu-\nu^{\prime})^{2}/\sigma^{2}]$, and $W_{p,k}(\nu)$ is a weight. Such weights are useful when some solution intervals contain fewer data points than others, due to interference excision, for example [^8] In this work, no interference is simulated and no such excision is performed, so these weights equal one..